Calculate the size of one of the interior angles of a regular decagon (i.e. a regular 10-sided polygon).
Enter the number of degrees to the nearest whole number in the box below.
(Your answer should be a whole number, without a degrees sign.)
-------------
I've tried to figure it out but it's to hard, are you able to help me please.
Thanks.

Answer by Alan3354(69443)

Calculate the size of one of the interior angles of a regular decagon (i.e. a regular 10-sided polygon).
--------------
The hard way:
Sum = 180*(n-2) = 180*8 = 1440 degs
1440/10 = 144 degs
---------------
Easy way:
Ext angle = 360/10 = 36 degs
Int angle = 180 - Ext = 144 degs